171 minus the first of the three consecutive integers is equivalent to the sum of the second and third. What are your integers?
I've tried several ways to work this problem but when I plug in the x, it doesn't make sense

Answer by ankor@dixie-net.com(22740)

The consecutive integers: x, (x+1), (x+2)
:
171 minus the first of the three consecutive integers is equivalent to the sum of the second and third.
171 - x = (x+1) + (x+2)
171 - x = 2x + 3
171 = 2x + x + 3
171 - 3 = 3x
168 = 3x
x = 168/3
x = 56
:
Try this



What are your integers?